# Mixed Planewave and Localized Orbital Basis for Sparse-Stochastic Hybrid TDDFT

### v1

@@ Abstract
We present a mixed basis-set approach to obtain optical absorption spectra within a generalized Kohn-Sham time-dependent density functional theory framework. All occupied valence molecular orbitals (MOs) are expanded in a plane-wave (PW) basis, while unoccupied MOs are derived primarily from localized atomic basis functions. The method accelerates spectral convergence when compared to fully PW-based simulations, with a $2-3$ fold reduction in the number of unoccupied MOs entering the Casida equation. The mixed-basis is placed on a common real-space grid, enabling our previously developed deterministic/sparse-stochastic evaluation of the exact exchange operator (DOI: 10.1021/acs.jctc.3c00987). This chemically intuitive and computationally efficient approach is validated across various molecular systems, including $\pi$-conjugated polymethine dyes, aromatic hydrocarbons, and a chlorophyll monomer.
preprint: Mixed Basis V3
I Introduction
Time-dependent density functional theory (TDDFT) is a widely used method for the excited-state properties of molecules, known for possessing a remarkable balance between accuracy and computational efficiency. An extension of ground-state DFT, TDDFT is built upon the Runge-Gross theorem which reduces the time evolution of the many-body electronic wavefunction to the independent propagation of single-particle molecular orbitals (MOs). Runge and Gross (1984) Since its inception, TDDFT has been used in predicting both linear and nonlinear optical response, photoelectron spectra, and optoelectronic properties in general.Casida (1995); Jamorski et al. (1996); Yabana and Bertsch (1997, 1999); Lopata et al. (2012); DeBeer George et al. (2008); Lermé et al. (1998); Qian et al. (2006); Baer and Neuhauser (2003); Baer and Gould (2001); Zelovich et al. (2014); Ding et al. (2013) Modern implementations of TDDFT employ hybrid functionals that incorporate a fraction of Fock exchange to the exchange-correlation (XC) potential. Inclusion of exact exchange provides a better description of delocalized excited states, such as $\pi\rightarrow\pi^{*}$ transitions, charge-transfer excitations, and excitations in systems with pronounced excitonic effects.Becke (1993); Baer et al. (2010); Leininger et al. (1997a); Yanai et al. (2004); Karolewski et al. (2011); Kronik et al. (2012); Vlček et al. (2019)
Going beyond the local density and generalized gradient approximations (LDA/GGA) Perdew and Wang (1992); Perdew et al. (1996) offers improved accuracy but adds to the computational cost with the need to evaluate the exchange integrals. In previous work, we introduced a mixed deterministic/sparse-stochastic exchange approach that reduces the scaling of both hybrid DFT and linear-response TDDFT within a plane-wave (PW) representation.Bradbury et al. (2023a); Sereda et al. (2024) This method splits the exchange interaction kernel, $u(k)$, into low- and high-$k$ components (where $k$ denotes a reciprocal lattice vector):
@@ I Introduction
In previous sparse-stochastic hybrid TDDFT calculations  Sereda et al. (2024), we found that many unoccupied MOs (roughly three to four times the number of occupied states) are required to obtain a converged absorption spectrum. This issue of convergence is well-known and inherent to the use of a PW basis-set. Booth et al. (2016); Sun et al. (2017) In this work, we resolve this issue by implementing a mixed plane-wave/atomic-orbital (PW/AO) basis-set representation for the MOs. The integration of mixed basis sets, combining PWs and AOs, is very useful in computational chemistry and materials science.Colle et al. (1987); Euwema (1971); Fabian et al. (2025); Cársky et al. (1998); Liu and Herbert (2015); Ko et al. (2021) For example, Samsonidze et al. demonstrated that unoccupied orbitals in GW calculations could be replaced with simple approximate physical orbitals, where Gaussian orbitals are used for resonant and continuum states.Samsonidze et al. (2011) Similarly, works conducted by Booth et al. and Sun et al. focus on Gaussian and PW hybrid representations alongside density fitting techniques.Booth et al. (2016); Sun et al. (2017)
@@ II Methodology
II.1 Iterative approach for optical absorption spectra
Excitation energies are obtained by solving the Casida equation Casida (1995):
@@ II.1 Iterative approach for optical absorption spectra
$\mathcal{L}=\begin{pmatrix}A&B\\ -B&-A\end{pmatrix},$ (3)
with $f^{+}$ and $f^{-}$ transition eigenvectors corresponding to excitations and de-excitations, respectively. For single-particle spin-singlet excitations, the $A$ and $B$ matrices read (assuming real-valued MOs throughout):
$A_{ia,jb}$ $=(\varepsilon_{a}-\varepsilon_{i})\delta_{ij}\delta_{ab}+2(ia|jb)$ (4)
@@
[…]
The matrix elements of the XC kernel, $f_{\text{XC}}$, are evaluated to first-order within an adiabatic LDA scheme. Baroni et al. (1987) The exchange integrals are calculated under a kernel $u$ that is split to short- and long-range parts in real space Becke (1993); Leininger et al. (1997b):
@@
Rather than direct diagonalization of Eq. (3), we use an iterative Chebyshev polynomial expansion approach to access the eigenvalues of $\mathcal{L}$, as in Refs. Bradbury et al. (2022, 2023b); Sereda et al. (2024). The absorption spectrum is calculated as:
@@
is a single exciton vector corresponding to a field polarization in the $\hat{r}$ direction. The Gaussian-broadened delta function is expanded in terms of Chebyshev polynomials:
$\delta(\mathcal{L}-\omega)=\sum_{n=0}^{N_{\text{Chebyshev}}}c_{n}(\omega)T_{n}(\tilde{\mathcal{L}}),$ (10)
where $T_{n}$ is the $n$’th-order Chebyshev polynomial and $\tilde{\mathcal{L}}$ is a scaled Liouvillian with eigenvalues between -1 and 1.Bradbury et al. (2023b); Weiße et al. (2006) $A(\omega)$ is therefore calculated from the Chebyshev moments,
@@
where $|\tilde{\chi}{\rangle}=\pm|\chi{\rangle}$ and for the coefficients, $c_{n}(\omega)$, simple smoothly decaying weights are used  Weiße et al. (2006).
II.2 Mixed-basis MOs on a real-space grid
@@
First, we perform two separate LDA-DFT calculations, one using a PW basis and the other with a localized AO basis. All vacuum molecular geometries are optimized in the ORCA 6.0 program at the B3LYP/def2-TZVPPD level of theory.Neese et al. (2020); Neese (2022) The PW calculations are performed using an in-house DFT code with norm-conserving pseudopotentials (NCPPs) representing the core electrons. Troullier and Martins (1991) The all-electron AO-basis calculations utilize the $PySCF$ package.Sun et al. (2018) The choice of basis is guided by benchmark calculations on the C20H10 molecule, where we compare the lowest-occupied MO (LUMO) - highest-occupied MO (HOMO) energy gaps obtained from LDA-DFT using various basis sets against a reference grid-based PW-LDA-DFT calculation. Among the tested basis-sets, aug-cc-pVDZ yields the smallest deviation from the PW result, as detailed in Appendix A.Dunning et al. (2001)
@@
In the PW-LDA-DFT calculations, the molecule is centered on a uniform real-space grid with a spacing of $dx=dy=dz=0.4$ Bohr and the Martyna-Tuckerman approach is used to minimize edge effects. Martyna and Tuckerman (1999). The simulation grids include at least 6 Bohr of padding beyond the extent of the molecule in all directions. Both the PW- and AO-MOs are then evaluated on this same grid. By replacing all or a subset of the unoccupied PW-MOs with those obtained from the AO calculation, we construct a mixed-basis representation of the MOs (Fig. 1).
@@
For static bandgaps we find that all PW-based unoccupied states can be replaced with a lesser number of AO-derived MOs and still reproduce the gaps of the fully-PW based calculations. However, a more careful selection of the virtual-space MOs is required for excited-state simulations. For sparse-stochastic hybrid TDDFT calculations, we observe that retaining every fourth unoccupied MO from the PW basis while replacing the rest with AO-derived orbitals provides the best agreement with the first excitation energies, i.e., absorption onset, obtained from fully PW-TDDFT. Specifically, the AO basis sets tend to overly localize, particularly in the virtual space, often resulting in a basis-set-induced blueshift. By retaining approximately 25% of the PW unoccupied MOs, we mitigate this shift and recover excitation energies consistent with fully PW-based calculations. To ensure a consistent set of MOs, we apply this same mixing strategy in both the static DFT and excited-state TDDFT simulations.
@@
Further, the mixing of an all-electron AO basis-set with an NCPP-PW approach results in lower HOMO energies in the PW-LDA-DFT calculations. To ensure energy alignment between the two frameworks, we apply a scissor-shift correction ($\Delta\varepsilon$), i.e., $\varepsilon_{i}^{PW}=\varepsilon_{i}^{AO}+\Delta\varepsilon$, where $\Delta\varepsilon$ is determined from the difference between the HOMO energies Samsonidze et al. (2011).
@@
II.3 Orthonormalization
Subsequently, we perform an overall orthonormalization of the mixed-basis MOs using QR decomposition by Householder transformation.Golub and Businger (1965); Press et al. (2007) This is done by decomposing the eigenvector matrix $\sqrt{dV}C$, where the columns of $C$ are non-orthogonal mixed LDA-DFT MOs and $dV$ is the volume element, into an orthonormal matrix $Q$ and an upper triangular matrix $R$. The columns of $Q$ correspond to the orthonormalized MOs on the grid. While other orthonormalization techniques, such as Gram-Schmidt or Löwdin transformation with singular value decomposition (SVD),Szabo and Ostlund (1996); Daniel et al. (1976); Golub and Reinsch (1970); Löwdin (1950); Bowler and Miyazaki (2012) are also applicable, we find the QR method via Householder transformation to be the most efficient. A Frobenius norm error analysis comparing different orthonormalization techniques is shown in Appendix B.
II.4 TDDFT Spectra and Spectral Convergence
With orthonormalized mixed-basis LDA MOs, we proceed with a sparse-stochastic hybrid DFT calculation using the CAM-LDA0 ($\alpha=0.19,\beta=0.46,\gamma=0.33$ $\rm{Bohr}^{-1}$).Yanai et al. (2004); Bradbury et al. (2023a); Sereda et al. (2024) This calculation updates the LDA-DFT Hamiltonian to include parametrized exact-exchange within a GKS-DFT framework. The hybrid exchange matrix elements are calculated using 1000 short stochastic fragments. For further details, see Ref. Bradbury et al. (2023a). Subsequently, we perform sparse-stochastic hybrid TDDFT calculations with 1000 Chebyshev terms to compute the optical absorption spectra. All occupied valence electron MOs are included in the valence space ($N_{v}=N_{\text{occ}}$), while in the TDDFT calculations the number of unoccupied MOs ($N_{c}$) is varied.
@@ II.4 TDDFT Spectra and Spectral Convergence
[…]
@@ III Results and Discussion
Figure 2: (a) HOMO, LUMO, and LUMO+90 densities and energies for C20H10 calculated at the LDA level with the mixed PW/AO@aug-cc-pVDZ basis (left) and pure PW basis (right). (b) MO energy comparison for C20H10 for unoccupied states between PW and AO@aug-cc-pVDZ basis at LDA level (HOMO energies have been aligned with a rigid energy-shift of $\Delta\varepsilon=-1.24$ eV). (c) Overlap intensity between unoccupied LDA MOs from PW and AO@aug-cc-pVDZ basis for C20H10.
@@
Table 1: Static LUMO-HOMO gaps (in eV) for four test systems from PW-LDA-DFT, AO@aug-cc-pVDZ LDA-DFT, PW- and mixed PW/AO- sparse-stochastic hybrid-DFT with CAM-LDA0.
@@
The overall convergence metrics $\delta$ and $\bar{\delta}$ are summarized in Table 2. The mixed-basis calculations show significantly lower $\delta$ and $\bar{\delta}$ values compared to their PW counterparts. Specifically, we observe a $2-3$ fold reduction in the values for $\delta$ and $\bar{\delta}$ using the mixed-based representation of the MOs relative to a purely PW-based representation. This enhanced performance originates from the more compact and chemically intuitive representation of high-energy conduction states in the AO basis, which effectively reduces the number of conduction states ($N_{c}$) required for achieving spectral convergence.
@@
However, our analysis also reveals an important subtlety regarding the first excitation energy, i.e., the optical gap. While the overall spectral shapes show excellent agreement between the PW and mixed-basis calculations, we consistently observe a small ($\sim$ 0.08 eV) but systematic blueshift in the optical gaps.Bruneval and Gonze (2008); Rocca et al. (2008) This observation suggests that, although the mixed basis is highly effective for describing higher-energy excitations and provides significantly faster spectral convergence, special care is required when precise determination of the optical gap is needed.
IV Conclusion
This study demonstrates the advantages of combining PW and AO basis sets within the sparse-stochastic hybrid TDDFT framework, offering a balanced approach that maintains accuracy while improving computational efficiency. The mixed-basis method, leveraging PW-derived occupied MOs and carefully selected AO-derived unoccupied MOs, achieves results with superior spectral convergence with respect to the number of conduction states. Our analysis confirms excellent agreement for ground-state properties in sparse-stochastic hybrid DFT calculations, where the LUMO-HOMO gap deviations between the PW and mixed-basis calculations are below 0.02 eV. The sparse-stochastic hybrid TDDFT spectral calculations show that the localized nature of AOs enhances convergence for high-energy excitations and offers an overall $2-3$ folds faster convergence.

### v2

@@ Abstract
We present a mixed basis-set approach to obtain optical absorption spectra within a generalized Kohn-Sham time-dependent density functional theory framework. All occupied valence molecular orbitals (MOs) are expanded in a plane-wave (PW) basis, while unoccupied MOs are derived primarily from localized atomic basis functions. The method accelerates spectral convergence when compared to fully PW-based simulations, with a 232-3 fold reduction in the number of unoccupied MOs entering the Casida equation. The mixed-basis is placed on a common real-space grid, enabling our previously developed deterministic/sparse-stochastic evaluation of the exact exchange operator (J. Chem. Theory Comput. 2023, 19, 9239–9247). This chemically intuitive and computationally efficient approach is validated across various molecular systems, including $\pi$-conjugated polymethine dyes, aromatic hydrocarbons, and a chlorophyll monomer.
[…]
These $\altaffiliationThese authors contributed equally.$ University of California, Los Angeles] Department of Chemistry and Biochemistry, University of California, Los Angeles, Los Angeles, California 90095, United States
{tocentry}
1 Introduction
Time-dependent density functional theory (TDDFT) is a widely used method for the excited-state properties of molecules, known for possessing a remarkable balance between accuracy and computational efficiency. An extension of ground-state DFT, TDDFT is built upon the Runge-Gross theorem which reduces the time evolution of the many-body electronic wavefunction to the independent propagation of single-particle molecular orbitals (MOs). 1 Since its inception, TDDFT has been used in predicting both linear and nonlinear optical response, photoelectron spectra, and optoelectronic properties in general.2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14 Modern implementations of TDDFT employ hybrid functionals that incorporate a fraction of Fock exchange to the exchange-correlation (XC) potential. Inclusion of exact exchange provides a better description of delocalized excited states, such as $\pi\rightarrow\pi^{*}$ transitions, charge-transfer excitations, and excitations in systems with pronounced excitonic effects.15, 16, 17, 18, 19, 20, 21
Going beyond the local density and generalized gradient approximations (LDA/GGA) 22, 23 offers improved accuracy but adds to the computational cost with the need to evaluate the exchange integrals. In previous work, we introduced a mixed deterministic/sparse-stochastic exchange approach that reduces the scaling of both hybrid DFT and linear-response TDDFT within a plane-wave (PW) representation.24, 25 This method splits the exchange interaction kernel, $u(k)$, into low- and high-$k$ components (where $k$ denotes a reciprocal lattice vector):
@@ 1 Introduction
In previous sparse-stochastic hybrid TDDFT calculations  25, we found that many unoccupied MOs (roughly three to four times the number of occupied states) are required to obtain a converged absorption spectrum. This issue of convergence is well-known and inherent to the use of a PW basis-set. 26, 27 In this work, we resolve this issue by implementing a mixed plane-wave/atomic-orbital (PW/AO) basis-set representation for the MOs. The integration of mixed basis sets, combining PWs and AOs, is very useful in computational chemistry and materials science.28, 29, 30, 31, 32, 33 For example, Samsonidze et al. demonstrated that unoccupied orbitals in GW calculations could be replaced with simple approximate physical orbitals, where Gaussian orbitals are used for resonant and continuum states.34 Similarly, works conducted by Booth et al. and Sun et al. focus on Gaussian and PW hybrid representations alongside density fitting techniques.26, 27
@@ 2 Methodology
2.1 A. Iterative approach for optical absorption spectra
Excitation energies are obtained by solving the Casida equation 2:
@@
$\mathcal{L}={\begin{pmatrix}\mathcal{A}&\mathcal{B}\\ -\mathcal{B}&-\mathcal{A}\end{pmatrix}}$, (3)
with $f^{+}$ and $f^{-}$ transition eigenvectors corresponding to excitations and de-excitations, respectively. For single-particle spin-singlet excitations, the $\mathcal{A}$} and $\mathcal{B}$} matrices read (assuming real-valued MOs throughout):
${\mathcal{A}_{ia,jb}}$ $=(\varepsilon_{a}-\varepsilon_{i})\delta_{ij}\delta_{ab}+2(ia|jb)$ (4)
@@
$(ia|jb)=\int dr\,dr^{\prime}\,\phi_{i}(r)\phi_{a}(r)|r-r^{\prime}|^{-1}\phi_{j}(r^{\prime})\phi_{b}(r^{\prime}).$ (6)
The matrix elements of the XC kernel, $f_{\text{XC}}$, are evaluated to first-order within an adiabatic LDA scheme. 35 The exchange integrals are calculated under a kernel $u$ that is split to short- and long-range parts in real space 15, 36:
@@
Rather than direct diagonalization of Eq. (3), we use an iterative Chebyshev polynomial expansion approach to access the eigenvalues of $\mathcal{L}$, as in Refs. 37, 38, 25, 12. The absorption spectrum is calculated as:
@@
is a single exciton vector corresponding to a field polarization in the $\hat{r}$ direction, combining forward (++) and backward (-) dipole transitions as detailed in Refs.12, 25 The Gaussian-broadened delta function is expanded in terms of Chebyshev polynomials:
[…]
where $T_{n}$ is the $n$’th-order Chebyshev polynomial and $\tilde{\mathcal{L}}$ is a scaled Liouvillian with eigenvalues between 1-1 and 11.38, 39 $A(\omega)$ is therefore calculated from the Chebyshev moments,
@@
where $|\tilde{\chi}{\rangle}=\pm|\chi{\rangle}$ and for the coefficients, $c_{n}(\omega)$, simple smoothly decaying weights are used.39
2.2 B. Mixed-basis MOs on a real-space grid
@@
First, we perform two separate LDA-DFT calculations, one using a PW basis and the other with a localized AO basis. All vacuum molecular geometries are optimized in the ORCA 6.0 program at the B3LYP/def2-TZVPPD level of theory.40, 41 The PW calculations are performed using an in-house DFT code with norm-conserving pseudopotentials (NCPPs) representing the core electrons. 42 The all-electron AO-basis calculations utilize the $PySCF$ package.43 The choice of basis is guided by benchmark calculations on the C20H10 molecule, where we compare the lowest-occupied MO (LUMO) - highest-occupied MO (HOMO) energy gaps obtained from LDA-DFT using various basis sets against a reference grid-based PW-LDA-DFT calculation. Among the tested basis-sets, aug-cc-pVDZ yields the smallest deviation from the PW result, as detailed in Appendix A.44
@@
In the PW-LDA-DFT calculations, the molecule is centered on a uniform real-space grid with a spacing of $dx=dy=dz=0.4$ Bohr. The simulation grids include at least 6 Bohr of padding beyond the extent of the molecule in all directions, this padding ensures proper boundary conditions and avoids periodic images in finite systems. The Martyna-Tuckerman approach is used to further minimize edge effects. 45. Both the PW- and AO-MOs are then evaluated on this same grid. By replacing all or a subset of the unoccupied PW-MOs with those obtained from the AO calculation, we construct a mixed-basis representation of the MOs (Fig. 1).
@@
For static bandgaps we find that all PW-based unoccupied states can be replaced with a lesser number of AO-derived MOs and still reproduce the gaps of the fully-PW based calculations. However, a more careful selection of the virtual-space MOs is required for excited-state simulations. For sparse-stochastic hybrid TDDFT calculations, we observe that retaining every fourth unoccupied MO from the PW basis while replacing the rest with AO-derived orbitals provides the best agreement with the first excitation energies, i.e., absorption onset, obtained from fully PW-TDDFT. Specifically, the AO basis sets tend to overly localize, particularly in the virtual space, often resulting in a basis-set-induced blueshift. By retaining approximately 25% of the PW unoccupied MOs, we mitigate this shift and recover excitation energies consistent with fully PW-based calculations. The 25% PW retaining ratio is optimized for molecular systems exhibiting naturally localized transitions, where our method demonstrates particular strength. To ensure a consistent set of MOs, we apply this same mixing strategy in both the static DFT and excited-state TDDFT simulations.
@@
Further, the mixing of an all-electron AO basis-set with an NCPP-PW approach results in lower HOMO energies in the PW-LDA-DFT calculations. To ensure energy alignment between the two frameworks, we apply a scissor-shift correction ($\Delta\varepsilon$), i.e., $\varepsilon_{i}^{PW}=\varepsilon_{i}^{AO}+\Delta\varepsilon$, where $\Delta\varepsilon$ is determined from the difference between the HOMO energies. 34 The HOMO energy alignment accounts for systematic differences between the all-electron AO and NCPP-PW treatments. Although the absolute energies differ due to their distinct representations of the electron-nuclear interaction, the preservation of energy gaps ensures the physical validity of the mixed-basis approach. The initial PW/AO orbital mixing occurs at the LDA level without self-consistency, requiring only HOMO alignment. We then use the mixed basis to perform the GKS-DFT SCF procedure.
@@
2.3 C. Orthonormalization
Subsequently, we perform an overall orthonormalization of the mixed-basis MOs using QR decomposition by Householder transformation.46, 47 This is done by decomposing the eigenvector matrix $\sqrt{dV}C$, where the columns of $C$ are non-orthogonal mixed LDA-DFT MOs and $dV$ is the volume element, into an orthonormal matrix $Q$ and an upper triangular matrix $R$. The columns of $Q$ correspond to the orthonormalized MOs on the grid. While other orthonormalization techniques, such as Gram-Schmidt or Löwdin transformation with singular value decomposition (SVD),48, 49, 50, 51, 52 are also applicable, we find the QR method via Householder transformation to be the most efficient. A Frobenius norm error analysis comparing different orthonormalization techniques is shown in Appendix B.
2.4 D. TDDFT Spectra and Spectral Convergence
With orthonormalized mixed-basis LDA MOs placed on a common real-space grid, we proceed with a sparse-stochastic hybrid DFT calculation using the CAM-LDA0 functional ($\alpha=0.19,\beta=0.46,\gamma=0.33$ $\rm{Bohr}^{-1}$).18, 24, 25 This calculation updates the LDA-DFT Hamiltonian to include parametrized exact-exchange within a GKS-DFT framework. The high-$k$ components of the hybrid exchange operator are represented with 1000 short stochastic fragments. The required number of stochastic fragments to obtain converged bandgaps and spectra (and the associated small stochastic error) has been extensively characterized in our previous work; for further details, see Ref. 24, 25. As the GKS-DFT calculation updates mixed-basis LDA MOs self-consistently, 1000 stochastic fragments uniformly sample both basis types while preserving their characters through the real-space representation. Subsequently, we perform sparse-stochastic hybrid TDDFT calculations with 1000 Chebyshev terms to compute the optical absorption spectra. All occupied valence electron MOs are included in the valence space ($N_{v}=N_{\text{occ}}$), while in the TDDFT calculations the number of unoccupied MOs ($N_{c}$) is varied. All TDDFT simulations use converged GKS-DFT MOs and eigenvalues within the mixed PW/AO representation as inputs into the Casida equation.
@@
$D_{ij}=\int|\tilde{A}_{i}(\omega)-\tilde{A}_{j}(\omega)|\;d\omega,$ (12)
where $\tilde{A}(\omega)$ is the area-normalized absorption spectrum, $A(\omega)$, i.e., $\tilde{A}(\omega)=A(\omega)/\int A(\omega)\,d\omega$. Consequently, $D_{ij}$ ranges between 0 (perfectly identical spectra) and 2 (completely different spectra). We further define two averaged quantities $\delta$ and $\bar{\delta}$,
@@ 3 Results and Discussion
Figure 2: (a) HOMO, LUMO, and LUMO+90 densities and energies for C20H10 calculated at the LDA level with the mixed PW/AO@aug-$c$-pVDZ basis (left) and pure PW basis (right). (b) MO energy comparison for C20H10 for unoccupied states between PW and AO@aug-$c$-pVDZ basis at LDA level (HOMO energies have been aligned with a rigid energy-shift of $\Delta\varepsilon=-1.24$ eV). (c) Overlap intensity between unoccupied LDA MOs from PW and AO@aug-$c$-pVDZ basis for C20H10.
@@
Table 1: Static LUMO-HOMO gaps (in eV) for four test systems from PW-LDA-DFT, AO@aug-$c$-pVDZ LDA-DFT, PW- and mixed PW/AO sparse-stochastic hybrid-DFT with CAM-LDA0 functional.
@@
The overall convergence metrics $\delta$ and $\bar{\delta}$ are summarized in Table 2. The mixed-basis calculations show significantly lower $\delta$ and $\bar{\delta}$ values compared to their PW counterparts. Specifically, we observe a 232-3 fold reduction in the values for $\delta$ and $\bar{\delta}$ using the mixed-based representation of the MOs relative to a purely PW-based representation. This enhanced performance originates from the more compact and chemically intuitive representation of high-energy conduction states in the AO basis, which effectively reduces the number of conduction states ($N_{c}$) required for achieving spectral convergence.
@@
We note that, while direct localization of PW virtual orbitals could offer an alternative, our mixed-basis approach ensures transferable accuracy without the need for state-specific tuning. However, our analysis also reveals an important subtlety regarding the first excitation energy, i.e., the optical gap. While the overall spectral shapes show excellent agreement between the PW and mixed-basis calculations, we consistently observe a small ($\sim$ 0.08 eV) but systematic blueshift in the optical gaps.53, 54 This observation suggests that, although the mixed basis is highly effective for describing higher-energy excitations and provides significantly faster spectral convergence, special care is required when precise determination of the optical gap is needed.
4 Conclusion
This study demonstrates the advantages of combining PW and AO basis sets within the sparse-stochastic hybrid TDDFT framework, offering a balanced approach that maintains accuracy while improving computational efficiency. The mixed-basis method, leveraging PW-derived occupied MOs and carefully selected AO-derived unoccupied MOs, achieves results with superior spectral convergence with respect to the number of conduction states. Our analysis confirms excellent agreement for ground-state properties in sparse-stochastic hybrid DFT calculations, where the LUMO-HOMO gap deviations between the PW and mixed-basis calculations are below 0.02 eV. The sparse-stochastic hybrid TDDFT spectral calculations show that the localized nature of AOs enhances convergence for high-energy excitations and offers an overall 232-3 folds faster convergence. We also acknowledge that while the current $\sim 25\%$ PW retaining ratio in the unoccupied manifold effectively mitigates blueshifts in neutral or positively charged organic systems, its extension to charged species or transition metal complexes will require further system-dependent adjustments of the mixing composition.